Homework 1, January 17

Math 402 – Spring Semester, 1990 Second Course in Abstract Algebra, II Prof. Daniel Grayson

2:00 PM MWF, Section F1, 243 Altgeld Hall

Read chapter IV: sections 1, 2, 4, 5, and 6; and chapter VII. Read the two chapters in parallel.

Do the exercises on p. 216: (1), (2), (3), 4, (7), (11), and on p. 354: (7), 8. The exercises in parentheses are to be turned in only if you are uncertain whether you got the correct answer. Turn them in Wednesday, January 24.